OpenStudy (anonymous):

Sasha's mom decided to get her a cell phone for her birthday. Cell Plus has a plan that costs $21.95 per month plus an additional $0.07 per minute. Cell Best has a plan that costs $25.95 per month plus an additional $0.05 per minute. How many minutes can Sasha talk and have the same cost each month? 0.12m = 4.00 21.95 + 0.07m = 26.00 21.95 + 0.07m = 25.95 + 0.05m 25.95 + 0.05m = 22.02

OpenStudy (valpey):

Do any of these seem like they correspond to the question?

OpenStudy (anonymous):

what do you mean?

OpenStudy (valpey):

Well, m is the variable referring to the number of minutes. In one plan the cost is a monthly charge plus 7 cents a min. In another plan it is a different monthly charge plus 5 cents a minute.

OpenStudy (valpey):

If Sasha only uses 10 minutes a month, clearly the cheaper monthly cost is more important. If she uses 10,000 minutes a month, clearly the lower per-minute cost is preferable.

OpenStudy (ganpat):

for Cell plus = 21.95 + 0.07x For Cell Best = 25.95 + 0.05 x where x is minutes.. Now for same cost, equate both.. 21.95 + 0.07x = 25.95 + 0.05 x 0.02x = 4 so x = 200 minutes...

OpenStudy (ganpat):

option C : 21.95 + 0.07m = 25.95 + 0.05m

OpenStudy (ganpat):

@rajfk35 , @Valpey : does it make sense ??